On a ECP Rectilinear Plant (1 DOF system) mass-spring-damper system, I am given the transfer function as: 1/(M*s^2 + B*s + K). With values mass M = 0.67538 kg, friction coefficient B = 1.8951 Ns/M, spring constant K = 322.278 N/M, and Damping coefficient d=2.54821 Ns/m.

I know the Open loop system model is: (1/(0.6738s^2 + 1.89515s + 322.278)) = (1.481/(s^2+2.806s+4.772))

Implement a simple controller using only one mass (+spring + damper) so the control is critically damped
Answer #1

1. For cvinca, ly damped Sy4km 4-1 Ta kirj N = 0.64538 KJ K= 322.23 8 시/m B2 = 4x,41538x 322.278 | B= 29.50 m=o-61538Kjk-322

Hope it helps you out.

In case any doubt please feel free to ask.

I will be happy to help you.

Thanks!!